OpenStudy (anonymous):

find the solution set: 2x + 7 \[\ge\] 3x - 1

OpenStudy (anonymous):

subtract 2x add 1

OpenStudy (anonymous):

\[2x+7\geq 3x-1\] \[7\geq x-1\] \[8\geq x\]

OpenStudy (anonymous):

thank you so much!! :))))

OpenStudy (anonymous):

yw

OpenStudy (nilankshi):

good